# Elucidated Rolling Diffusion Models for Probabilistic Weather Forecasting

### v1

@@ D.3 Baseline details
We follow the default values used by DYffusion for its Navier-Stokes experiments, but replace the interpolator and forecaster architecture with the more performant U-Net and use our optimization hyperparameters, both described above. We tuned the interpolator dropout rates and found a rate of 0.5 to produce the best results. The interpolator dropout rate is a key hyperparameter in DYffusion, since it is responsible for the stochasticity in the model (if 0, it reduces to a deterministic model). As for ERDM, we use a window size of $W=6$. Our improved architecture and training hyperparameters jointly improved the DYffusion CRPS scores by over 3×3$\times$3 × compared to the results in the original paper.
@@
Our primary focus is to evaluate ERDM’s performance relative to common EDM-derived approaches for dynamics forecasting. Consequently, our key baseline is an EDM denoiser parameterized by $D_{\theta}({\bm{x}}_{1}\ldots,{\bm{x}}_{W};\sigma,{\bm{y}}_{0})$, where all ${\bm{x}}_{w}$ are corrupted according to the same noise standard deviation, $\sigma$. With $W=1$, we recover a next-step forecasting conditional EDM as in [42, 26]. On Navier-Stokes, we tried $W\in\{1,2,4,6\}$, and found $W=4$ to work best for EDM. For computational reasons, we only report an EDM run with $W=1$ for ERA5. To condition the denoiser network on the initial condition, ${\bm{y}}_{0}$, we concatenate it along the channel dimension to the input, noisy window, ${\bm{x}}_{1:W}$. When $W>1$, ${\bm{y}}_{0}$ is thus first duplicated along the window dimension. We found this to perform better than operating on a $W+1$ window, where ${\bm{y}}_{0}$ is inserted into the first position of the window. Preconditioning is applied before concatenation, only to the noisy window.
@@ D.4 Distinctions between GenCast and our EDM model for ERA5
Our primary baseline, conditional EDM, applied to ERA5 shares several similarities with GenCast [42]. A primary distinction lies in the spatial resolution: GenCast is trained on a spatial resolution of $1^{\circ}$ and $0.25^{\circ}$, whereas our model, EDM, and other baselines are exclusively trained at a $1.5^{\circ}$ resolution. The denoiser network architecture presents another significant divergence. GenCast uses a graph neural network, similar to the approach in Lam et al. [27], while we use a 2D U-Net for our EDM baseline borrowed from [23, 12]. Furthermore, GenCast incorporates SST as both an input and an output variable. sea surface temperature as an input and output, which we do not. More minor differences between GenCast vs. our EDM are: ODE solver (DPMSolver++ vs. Heun), training noise level distribution (uniform vs. lognormal), noise distribution (spherical vs. i.i.d.), optimization hyperparameters, and residual vs. full-state prediction.
@@ D.5 Compute resources
For Navier-Stokes experiments, we trained all models and baselines on $2-8$ L40S or A100 GPUs. For ERA5, we trained some models (including our final ERDM version) on 4 H200 GPUs, but performed most of our development work on $8-16$ A100 GPUs (per training run). For ERDM with $W=6$ ($W=4$), a full training run on 8 L40S GPUs took around 16 (10.5) hours. For ERA5 with $W=6$, a full training run on 4 H200 GPUs took almost 5 days.
@@ D.6 Metrics
Let ${\bm{y}}\in\mathbb{R}^{I\times J}$ indicate the targets for a specific time step, and $\hat{\bm{y}}\in\mathbb{R}^{M\times I\times J}$ the corresponding predictions with an ensemble size of $M$. $I$ and $J$ are the size of the height dimension and width dimension, respectively. For ERA5, these are latitude and longitude, respectively. The definitions of the metrics below follow common practices (e.g., cf. Weatherbench-2 [43]).
@@
For ERA5 experiments, we follow the common practice of area-weighting all spatially aggregated metrics according to the size of the grid cell, ensuring that they are not biased towards the polar regions [43]. The unnormalized area weights are computed as $\tilde{w}(i)=\sin\phi_{i}^{u}-\sin\phi_{i}^{l}$, where $\phi_{i}^{u}$ and $\phi_{i}^{l}$ represent upper and lower latitude bounds, respectively, for the grid cell with latitude index $i\in\{1,2,\dots,I\}$. The normalized area weights, used in all metrics below are thus $w(i)=\frac{\tilde{w}(i)}{\frac{1}{I}\sum_{i=1}^{I}\tilde{w}(i)}$, where $I$ is the number of latitude indices. For the Navier-Stokes experiments, we report unweighted spatially aggregated metrics (i.e., $w(i)=1$ for all $i$).
@@
Following Fortin et al. [14], the spread-skill ratio is defined as the ratio between the ensemble spread and the ensemble-mean RMSE. The ensemble spread is defined as the square root of the ensemble variance. Thus, the SSR is defined as:
@@
Following [55, 43], we use the unbiased version of the CRPS [34] (lower is better)
@@
The CRPSS is commonly used to compare the CRPS scores relative to a baseline. It is computed as $1-\mathrm{CRPS}_{\text{model}}/\mathrm{CRPS}_{\text{baseline}}$, where $\mathrm{CRPS}_{\text{model}}$ and $\mathrm{CRPS}_{\text{baseline}}$ refer to the CRPS scores for the evaluated model and baseline, respectively. Values larger than 00 indicate that the model is better than the baseline model in underscore, while values smaller than 0 indicate that the model is worse than the baseline model.
@@ Appendix E Ablations
Starting from our proposed ERDM configuration (“Base” in Table 1), which uses the second-order Heun sampler, loss weighting with $P_{\text{mean}}=0.5,P_{\text{std}}=1.2$, a temporally correlated progressive noise prior following [15] with $\epsilon_{\rm prog}:=\alpha=1$ (see Section C.4), noise schedule parameters $\sigma_{\text{min}}=0.002,\sigma_{\text{max}}=200,\rho=-10$, and a window size of $W=6$, we systematically varied individual components. For sampling, we use $N=1.25$ ($\Delta t=0.8$) and $S_{\text{churn}}=0$ by default. The first window is initialized based on predictions from the EDM $W=4$ baseline. Our findings, summarized in Table 1, highlight several critical elements for achieving optimal performance.
Table 1: Ablation study results on the Navier-Stokes dataset. The most important design choices, without which ERDM would degrade to worse performance than the EDM baseline (highlighted in red), are 1) a proper 3D architecture; 2) an appropriate noise schedule (in particular, not EDM’s default $\rho=7$); 3) our proposed loss weighting; 4) a fixed training schedule. CRPSS refers to the CRPS skill score, where $>0$ means that the model is better than the baseline model in underscore. CRPSS values smaller than 0 indicate that the model is worse than the baseline model. For all other metrics, lower is better.
@@
Compared to the first-order Euler method (Algorithm 2), the second-order sampling algorithm detailed in Algorithm 3 achieved superior performance in CRPS, MSE, and SSR metrics. This improvement comes at the cost of doubling the neural function evaluations due to the correction step, which significantly reduces sampling speed. Nevertheless, we deemed this increased computational overhead justifiable for the enhanced accuracy.
@@
As detailed in the main text, a well-tuned loss weighting can significantly improve results. In particular, using no loss reweighting based on the PDF of the lognormal distribution, $f(\sigma;P_{\text{mean}},P_{\text{std}})$ (“No $f(\sigma)$”), degraded CRPS and MSE scores by more than 2×2$\times$2 × and 5×5$\times$5 ×, respectively. It also resulted in a worse calibration score. Similarly, choosing a poor value for the center of the lognormal distribution, e.g., $P_{\text{mean}}=-1.2$, resulted in almost as suboptimal results. Interestingly, $P_{\text{mean}}=-1.2$ corresponds to the default value in EDM, but we found values $\geq 0.5$ to be necessary for ERDM. We found EDM’s default $P_{\text{std}}=1.2$ to work well for ERDM generally.
@@
As hinted in Section C.4, we found it beneficial to use a temporally correlated noise prior as opposed to simply sampling i.i.d. noise for each snapshot (“$\epsilon_{\mathrm{indep}}$”). The noise prior that we use [15] introduces one hyperparameter, $\alpha$, which balances the strength of the correlation. We generally found $\alpha\in[0.5,1.5]$ to work best. Similarly to the findings of Ge et al. [15], using overly large $\alpha$ (e.g., $\alpha=10$, which introduces large correlations between noise snapshots) is detrimental to performance and worse than simply using i.i.d. noise.
@@ E.1 ERDM initialization
In Fig. 7 and the last part of Table 1, we ablate possible initialization schemes for ERDM. These include two neural initializations based on the EDM baselines with $W=1$ and $W=4$ (our default for Navier-Stokes being EDM $W=4$), alongside two simple non-neural approaches. One non-neural method, Init=truth, initializes ERDM using ground truth data for the first window. While clearly impractical, it serves as an informative upper bound on the achievable performance of ERDM. The other, Init=persistence, populates the first window by duplicating the initial condition ${\bm{y}}_{0}$; this approach performed very poorly and is omitted from Fig. 7 for visual clarity. Expectedly, Init=truth yields optimal initial CRPS results but cannot be used in practice. More significantly, Fig. 7 demonstrates that the choice of initialization predominantly impacts ERDM’s short-range forecast skill (up to approximately timestep 10). The initial performance advantages conferred by methods like Init=truth, or differences among other practical schemes, diminish substantially as the forecast horizon extends. By the end of the rollouts, both CRPS and SSR scores become largely indistinguishable across the various practical initialization methods, including our default EDM $W=4$ initialization. This convergence suggests that while the first-window initialization is crucial for early forecast accuracy, its specific choice becomes less critical for the long-range performance of ERDM (except when using a naive initialization such as persistence).
@@ E.2 ERA5 evaluation year ablation
In our quantitative comparison, we evaluated ERDM, EDM, and IFS ENS on the same 64 initial conditions evenly spaced in 2021, ensuring a fair comparison. Unfortunately, we only have access to the NeuralGCM ENS and Graph-EFM results for 2020. This discrepancy highlights the importance of understanding how model scores might vary across different evaluation years. We investigate this for our ERDM model (trained on data up to 2020 inclusive) in Fig. 8. As detailed in the figure and its caption, ERDM’s CRPS performance indeed shows sensitivity to the evaluation year. For example, when evaluated on 2022 initial conditions, scores often improved relative to our 2021 baseline, though for most variables this change was less than $3\%$. However, certain variables, notably geopotential (with z500 improving by up to $15\%$) and mean sea level pressure, exhibited significantly greater sensitivity, particularly at lead times under 8 days. These findings suggest that while our 2021-based evaluations provide a consistent benchmark for ERDM, EDM, and IFS ENS, direct comparisons with results from NeuralGCM ENS and Graph-EFM on 2020 data should be made with an awareness of this potential inter-annual performance variability.
@@ F.1 Additional Navier-Stokes results
For completeness, we report the ensemble-mean MSE test rollout scores for all Navier-Stokes models in Fig. 9, analogously to our main CRPS and SSR results in Fig. 4. The MSE scores are strongly correlated to the CRPS scores, which is why we deferred their discussion to the appendix. Similarly to the CRPS scores, ERDM achieves significantly better MSE scores, especially for long-range lead times. A slight difference to the CRPS results is that the relative improvement of the block-wise autoregressive EDM $W=4$ baseline over the DYffusion baseline is larger in terms of MSE than CRPS.
Generated videos.
Generated videos, the corresponding ground truth, and their lead time-dependent errors are provided at this URL: https://drive.google.com/drive/folders/1lVp9TwiK3dcQtsvr5XKlWiBEp1nidnum?usp=sharing.
@@ F.2 Detailed ERA5 quantitative results
Figure 10: Absolute CRPS corresponding to the relative scores shown in Fig. 5. Lower is better.
[…]
@@
A comprehensive quantitative assessment of forecast skill is presented in Fig. 11, which displays the CRPS for ERDM, IFS ENS, NeuralGCM, and Graph-EFM, relative to the EDM baseline. These relative CRPS values are reported as percentages (where lower indicates better performance) for numerous atmospheric variables (extending the set shown in Fig. 5) over the full 15-day forecast horizon. These results underscore that ERDM consistently achieves significantly lower CRPS values than the EDM baseline. The corresponding absolute CRPS scores are visualized in Fig. 10. A full relative comparison between ERDM and EDM is presented in the form of a scorecard in Fig. 13. A quantitative assessment of forecast calibration is presented in  Fig. 12, which displays the spread-skill ratio for the same set of models and variables. ERDM generally generates well-calibrated forecasts on par or better than IFS ENS, and consistently more calibrated forecasts than the EDM baseline or NeuralGCM.
[…]

### v3

@@ D.3 Baseline details
We follow the default values used by DYffusion for its Navier-Stokes experiments, but replace the interpolator and forecaster architecture with the more performant U-Net and use our optimization hyperparameters, both described above. We tuned the interpolator dropout rates and found a rate of 0.5 to produce the best results. The interpolator dropout rate is a key hyperparameter in DYffusion, since it is responsible for the stochasticity in the model (if 0, it reduces to a deterministic model). As for ERDM, we use a window size of $W=6$. Our improved architecture and training hyperparameters jointly improved the DYffusion CRPS scores by over $3\times$ compared to the results in the original paper.
PDE-Refiner
PDE-Refiner [lippe2023pderefiner] is a next-step diffusion-derived forecaster. We tune the two key hyperparameters of PDE-Refiner, $K$ and $\sigma_{\text{min}}$. The best forecasts were obtained with $K=6,\sigma_{\text{min}}=0.002$. The recommended values from [lippe2023pderefiner], $K=4,\sigma_{\text{min}}=2e\text{-}7$, did not work as well. We use the same denoiser architecture as for EDM. PDE-Refiner performs extremely well for the first few timesteps of the Navier-Stokes rollout, but rapidly diverges after that. We believe this is because it cannot model temporal interactions.
@@
Our primary focus is to evaluate ERDM’s performance relative to common EDM-derived approaches for dynamics forecasting. Consequently, our key baseline is an EDM denoiser parameterized by $D_{\theta}({\bm{x}}_{1}\ldots,{\bm{x}}_{W};\sigma,{\bm{y}}_{0})$, where all ${\bm{x}}_{w}$ are corrupted according to the same noise standard deviation, $\sigma$. With $W=1$, we recover a next-step forecasting conditional EDM as in [price2023gencast, kohl2023benchmarking]. On Navier-Stokes, we tried $W\in\{1,2,4,6\}$, and found $W=4$ to work best for EDM. For computational reasons, we only report an EDM run with $W=1$ for ERA5. To condition the denoiser network on the initial condition, ${\bm{y}}_{0}$, we concatenate it along the channel dimension to the input, noisy window, ${\bm{x}}_{1:W}$. When $W>1$, ${\bm{y}}_{0}$ is thus first duplicated along the window dimension. We found this to perform better than operating on a $W+1$ window, where ${\bm{y}}_{0}$ is inserted into the first position of the window. Preconditioning is applied before concatenation, only to the noisy window.
To be clear, the EDM “video” baseline with $W>1$ does not use rolling diffusion. It is a standard sequence-to-sequence EDM, using the same 3D denoiser architecture as ERDM, that operates autoregressively. Conditioned on a single clean frame (e.g., at time $t$), it jointly predicts a window of future frames ($t+1$ to $t+W$). Unlike ERDM, all frames in this window are denoised from the same initial noise level. To generate a long forecast, the last frame of the predicted window (at time $t+W$) is then used as the new “clean” condition for the next prediction step.
@@ D.4 Distinctions between GenCast and our EDM model for ERA5
Our primary baseline, conditional EDM, applied to ERA5 shares several similarities with GenCast [price2023gencast]. A primary distinction lies in the spatial resolution: GenCast is trained on a spatial resolution of $1^{\circ}$ and $0.25^{\circ}$, whereas our model, EDM, and other baselines are exclusively trained at a $1.5^{\circ}$ resolution. The denoiser network architecture presents another significant divergence. GenCast uses a graph neural network, similar to the approach in lam2023learning, while we use a 2D U-Net for our EDM baseline borrowed from [karras2022edm, dhariwal2021diffusion]. Furthermore, GenCast incorporates SST as both an input and an output variable. sea surface temperature as an input and output, which we do not. More minor differences between GenCast vs. our EDM are: ODE solver (DPMSolver++ vs. Heun), training noise level distribution (uniform vs. lognormal), noise distribution (spherical vs. i.i.d.), optimization hyperparameters, and residual vs. full-state prediction.
@@ D.5 Compute resources
For Navier-Stokes experiments, we trained all models and baselines on 282-8 L40S or A100 GPUs. For ERA5, we trained some models (including our final ERDM version) on 4 H200 GPUs, but performed most of our development work on 8168-16 A100 GPUs (per training run). For ERDM with $W=6$ ($W=4$), a full training run on 8 L40S GPUs took around 16 (10.5) hours. For ERA5 with $W=6$, a full training run on 4 H200 GPUs took almost 5 days.
@@ D.6 Metrics
Let ${\bm{y}}\in\mathbb{R}^{I\times J}$ indicate the targets for a specific time step, and $\hat{\bm{y}}\in\mathbb{R}^{M\times I\times J}$ the corresponding predictions with an ensemble size of $M$. II and $J$ are the size of the height dimension and width dimension, respectively. For ERA5, these are latitude and longitude, respectively. The definitions of the metrics below follow common practices (e.g., cf. Weatherbench-2 [rasp2023weatherbench2]).
@@
For ERA5 experiments, we follow the common practice of area-weighting all spatially aggregated metrics according to the size of the grid cell, ensuring that they are not biased towards the polar regions [rasp2023weatherbench2]. The unnormalized area weights are computed as $\tilde{w}(i)=\sin\phi_{i}^{u}-\sin\phi_{i}^{l}$, where $\phi_{i}^{u}$ and $\phi_{i}^{l}$ represent upper and lower latitude bounds, respectively, for the grid cell with latitude index $i\in\{1,2,\dots,I\}$. The normalized area weights, used in all metrics below are thus $w(i)=\frac{\tilde{w}(i)}{\frac{1}{I}\sum_{i=1}^{I}\tilde{w}(i)}$, where $I$ is the number of latitude indices. For the Navier-Stokes experiments, we report unweighted spatially aggregated metrics (i.e., $w(i)=1$ for all $i$).
@@
Following fortin2014ssr, the spread-skill ratio is defined as the ratio between the ensemble spread and the ensemble-mean RMSE. The ensemble spread is defined as the square root of the ensemble variance. Thus, the SSR is defined as:
@@
Following [zamo2018faircrps, rasp2023weatherbench2], we use the unbiased version of the CRPS [matheson1976crps] (lower is better)
@@
The CRPSS is commonly used to compare the CRPS scores relative to a baseline. It is computed as $1-\mathrm{CRPS}_{\text{model}}/\mathrm{CRPS}_{\text{baseline}}$, where $\mathrm{CRPS}_{\text{model}}$ and $\mathrm{CRPS}_{\text{baseline}}$ refer to the CRPS scores for the evaluated model and baseline, respectively. Values larger than 0 indicate that the model is better than the baseline model in underscore, while values smaller than 0 indicate that the model is worse than the baseline model.
@@ Appendix E Ablations
Starting from our proposed ERDM configuration (“Base” in Table˜2), which uses the second-order Heun sampler, loss weighting with $P_{\text{mean}}=0.5,P_{\text{std}}=1.2$, a temporally correlated progressive noise prior following [ge2024noiseprior] with $\epsilon_{\rm prog}:=\alpha=1$ (see Section˜C.4), noise schedule parameters $\sigma_{\text{min}}=0.002,\sigma_{\text{max}}=200,\rho=-10$, and a window size of $W=6$, we systematically varied individual components. For sampling, we use $N=1.25$ ($\Delta t=0.8$) and $S_{\text{churn}}=0$ by default. The first window is initialized based on predictions from the EDM $W=4$ baseline. Our findings, summarized in Table˜2, highlight several critical elements for achieving optimal performance.
Table 2: Ablation study results on the Navier-Stokes dataset. The most important design choices, without which ERDM would degrade to worse performance than the EDM baseline (highlighted in red), are 1) a proper 3D architecture; 2) an appropriate noise schedule (in particular, not EDM’s default $\rho=7$); 3) our proposed loss weighting; 4) a fixed training schedule. CRPSS refers to the CRPS skill score, where >0>0 means that the model is better than the baseline model in underscore. CRPSS values smaller than 0 indicate that the model is worse than the baseline model. For all other metrics, lower is better.
@@
Compared to the first-order Euler method (Algorithm˜2), the second-order sampling algorithm detailed in Algorithm˜3 achieved superior performance in CRPS, MSE, and SSR metrics. This improvement comes at the cost of doubling the neural function evaluations due to the correction step, which significantly reduces sampling speed. Nevertheless, we deemed this increased computational overhead justifiable for the enhanced accuracy.
@@
As detailed in the main text, a well-tuned loss weighting can significantly improve results. In particular, using no loss reweighting based on the PDF of the lognormal distribution, $f(\sigma;P_{\text{mean}},P_{\text{std}})$ (“No $f(\sigma)$”), degraded CRPS and MSE scores by more than $2\times$ and $5\times$, respectively. It also resulted in a worse calibration score. Similarly, choosing a poor value for the center of the lognormal distribution, e.g., $P_{\text{mean}}=-1.2$, resulted in almost as suboptimal results. Interestingly, $P_{\text{mean}}=-1.2$ corresponds to the default value in EDM, but we found values $\geq 0.5$ to be necessary for ERDM. We found EDM’s default $P_{\text{std}}=1.2$ to work well for ERDM generally.
@@
As hinted in Section˜C.4, we found it beneficial to use a temporally correlated noise prior as opposed to simply sampling i.i.d. noise for each snapshot ($\epsilon_{\mathrm{indep}}$”). The noise prior that we use [ge2024noiseprior] introduces one hyperparameter, $\alpha$, which balances the strength of the correlation. We generally found $\alpha\in[0.5,1.5]$ to work best. Similarly to the findings of ge2024noiseprior, using overly large $\alpha$ (e.g., $\alpha=10$, which introduces large correlations between noise snapshots) is detrimental to performance and worse than simply using i.i.d. noise.
@@ E.1 ERDM initialization
In Fig.˜7 and the last part of Table˜2, we ablate possible initialization schemes for ERDM. These include two neural initializations based on the EDM baselines with $W=1$ and $W=4$ (our default for Navier-Stokes being EDM $W=4$), alongside two simple non-neural approaches. One non-neural method, Init=truth, initializes ERDM using ground truth data for the first window. While clearly impractical, it serves as an informative upper bound on the achievable performance of ERDM. The other, Init=persistence, populates the first window by duplicating the initial condition ${\bm{y}}_{0}$; this approach performed very poorly and is omitted from Fig.˜7 for visual clarity. Expectedly, Init=truth yields optimal initial CRPS results but cannot be used in practice. More significantly, Fig.˜7 demonstrates that the choice of initialization predominantly impacts ERDM’s short-range forecast skill (up to approximately timestep 10). The initial performance advantages conferred by methods like Init=truth, or differences among other practical schemes, diminish substantially as the forecast horizon extends. By the end of the rollouts, both CRPS and SSR scores become largely indistinguishable across the various practical initialization methods, including our default EDM $W=4$ initialization. This convergence suggests that while the first-window initialization is crucial for early forecast accuracy, its specific choice becomes less critical for the long-range performance of ERDM (except when using a naive initialization such as persistence).
@@ E.2 ERA5 evaluation year ablation
In our quantitative comparison, we evaluated ERDM, EDM, and IFS ENS on the same 64 initial conditions evenly spaced in 2021, ensuring a fair comparison. Unfortunately, we only have access to the NeuralGCM ENS and Graph-EFM results for 2020. This discrepancy highlights the importance of understanding how model scores might vary across different evaluation years. We investigate this for our ERDM model (trained on data up to 2020 inclusive) in Fig.˜8. As detailed in the figure and its caption, ERDM’s CRPS performance indeed shows sensitivity to the evaluation year. For example, when evaluated on 2022 initial conditions, scores often improved relative to our 2021 baseline, though for most variables this change was less than $3\%$. However, certain variables, notably geopotential (with z500 improving by up to $15\%$) and mean sea level pressure, exhibited significantly greater sensitivity, particularly at lead times under 8 days. These findings suggest that while our 2021-based evaluations provide a consistent benchmark for ERDM, EDM, and IFS ENS, direct comparisons with results from NeuralGCM ENS and Graph-EFM on 2020 data should be made with an awareness of this potential inter-annual performance variability.
@@ F.1 Additional Navier-Stokes results
For completeness, we report the ensemble-mean MSE test rollout scores for all Navier-Stokes models in Fig.˜9, analogously to our main CRPS and SSR results in Fig.˜4. The MSE scores are strongly correlated to the CRPS scores, which is why we deferred their discussion to the appendix. Similarly to the CRPS scores, ERDM achieves significantly better MSE scores, especially for long-range lead times. A slight difference to the CRPS results is that the relative improvement of the block-wise autoregressive EDM $W=4$ baseline over the DYffusion baseline is larger in terms of MSE than CRPS.
@@ F.2 Detailed ERA5 quantitative results
Figure 10: Absolute CRPS corresponding to the relative scores shown in Fig.˜5. Lower is better.
Figure 11: CRPS relative to next-step autoregressive EDM baseline (in $\%$; lower is better) for 15-day rollouts across a more comprehensive set of variables than in Fig.˜5.
@@
A comprehensive quantitative assessment of forecast skill is presented in Fig.˜11, which displays the CRPS for ERDM, IFS ENS, NeuralGCM, and Graph-EFM, relative to the EDM baseline. These relative CRPS values are reported as percentages (where lower indicates better performance) for numerous atmospheric variables (extending the set shown in Fig.˜5) over the full 15-day forecast horizon. These results underscore that ERDM consistently achieves significantly lower CRPS values than the EDM baseline. The corresponding absolute CRPS scores are visualized in Fig.˜10, and a selection of them are summarized in Table˜3. A full relative comparison between ERDM and EDM is presented in the form of a scorecard in Fig.˜13. A quantitative assessment of forecast calibration is presented in  Fig.˜12, which displays the spread-skill ratio for the same set of models and variables. ERDM generally generates well-calibrated forecasts on par or better than IFS ENS, and consistently more calibrated forecasts than the EDM baseline or NeuralGCM.
Table 3: Tabular CRPS scores for a selection of forecast horizons (three, seven, and ten days) and variables (10m u-component of wind and 500 hPa geopotential). Lower is better.
u10m z500
Model 3d 7d 14d 3d 7d 14d
EDM 0.6923 1.324 1.769 76.04 204.9 327.4
ERDM 0.6748 1.312 1.737 69.29 197.5 316.7
IFS ENS 0.6830 1.318 1.785 58.41 183.4 317.3
NeuralGCM N/A 54.59 173.1 311.1
Figure 13: Scorecard of ERDM’s CRPS relative to next-step autoregressive EDM baseline (in $\%$; lower is better) for all predicted variables (except v, which is correlated with u), pressure levels, and lead times (up to 15 days). ERDM consistently outperforms the EDM baseline, especially for geopotential and high-altitude levels.